OpenStudy (anonymous):

i know im going to sound a little stupid but im having issues with my little sister's 2nd grade homework. Which is a doubles minus one fact? - 7+7=14 - 7+5=12 - 8+7=15 - 8+8=16 With explanation please because I want to be able to teach her!

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (immanuelv):

https://www.youtube.com/watch?v=elj4aup0wJk

OpenStudy (loser66):

do you mean -(-7)+7 =14?

OpenStudy (anonymous):

no thats what it shows me @Loser66

OpenStudy (compassionate):

"Which is a doubles minus one fact?"

OpenStudy (loser66):

if it is -7+7 , then it is =0 , cannot be 14 if it is -(-7) +7 , then it is = 14

OpenStudy (anonymous):

there is two minus one facts though @Compassionate

OpenStudy (anonymous):

@Loser66 no what i typed in is what the paper shows me

OpenStudy (loser66):

then, It's above my head. I am sorry for being helpless :)

OpenStudy (anonymous):

a) 7+7=14 b) 7+5=12 c) 8+7=15 d) 8+8=16 @Loser66

OpenStudy (anonymous):

its alright :)

OpenStudy (anonymous):

im literally getting no help

OpenStudy (anonymous):

Double facts are relatively easy to learn and become powerful anchors for other addition facts. Students quickly learn and identify patterns in the below sequence. 1 + 1 = 2 6 + 6 = 12 2 + 2 = 4 7 + 7 = 14 3 + 3 = 6 8 + 8 = 16 4 + 4 = 8 9 + 9 = 18 5 + 5 = 10 10 + 10 = 20

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

its doubles minus one fact so not really :) @woods33

OpenStudy (anonymous):

Doubles Minus one is very similar, except you take the larger addend and double it. This time you will subtract one. Example 7+6= 7+7=14-1=13

OpenStudy (anonymous):

its telling her to choose one

OpenStudy (anonymous):

that help out

OpenStudy (anonymous):

so the answer 7+7=14 ?

OpenStudy (anonymous):

yes 7+7=14